# Autonomy Architectures for Safe Planning in Unknown Environments Under Budget Constraints

### v1

@@ III-A Guaranteed Constraint Satisfaction
Then since $p_{k+1}^{\textup{com}}$ is dynamically feasible by construction, $x(t)=p_{k+1}^{\textup{com}}(t)\quad\forall t\in[t_{k+1},t_{k+2})$. Therefore, $x(t)=p_{k}^{\textup{com}}(t)\in\mathcal{S}~{}\forall t\in[t_{k},t_{k+1}),~{}\forall k\in\mathbb{N}\implies x(t)\in\mathcal{S}~{}\forall t\geq t_{0}$. Moreover, the solution to (3) with the corresponding $x(t)$ and $u(t)$ satisfies $b(t)\leq B~{}\forall t\geq t_{0}$. ∎
We have shown that using the gatekeeper architecture with a backup trajectory that reaches $\mathcal{R}_{k}$, the mission requirements are satisfied for all time. Next, we introduce a method for constructing backup trajectories and selecting one that minimizes the budget state $b$.
III-B Backup Trajectory Construction
Our method relies on posing a trajectory optimization problem and solving it with a sampling-based algorithm called ReRoot.
III-B1 Trajectory Optimization Problem
@@
where $t_{kB}=t_{k}+T_{S}+T_{B}$ is the finite time when $p$ reaches $\mathcal{R}_{k}$.
For any switch time $T_{S}$, the ideal backup trajectory is the one that minimizes the budget expenditure. Given a known budget renewal set $\mathcal{R}_{k}$ at time $t_{k}$, design a backup trajectory planner that computes a dynamically feasible trajectory $p^{\textup{bck},T_{S}}_{k}$ starting at $p_{k}^{\textup{can},T_{S}}(t_{kS})$ and ending in $\mathcal{R}_{k}$ while remaining in the known safe set $\mathcal{F}_{k}$ with minimal budget expenditure along the backup trajectory:
$\min_{\begin{subarray}{c}p(\cdot),b(\cdot),u(\cdot)\end{subarray}}\quad$ $b^{-}$ (22a)
s.t. $\dot{p}(t)=f(x(t),u(t))$ (22b)
@@
$p(t_{kS})=p_{k}^{\textup{can},T_{S}}(t_{kS})$ (22d)
$b(t_{kS})=b_{k}^{\textup{can},T_{S}}(t_{kS})$ (22e)
$x(t)\in\mathcal{F}_{k}\quad\forall t\in[t_{kS},t_{kB}]$ (22f)
$u(t)\in\mathcal{U}\quad\forall t\in[t_{kS},t_{kB}]$ (22g)
$p(t_{kB})\in\mathcal{R}_{k}$ (22h)
Here, (22b) and (22c) are the robot and budget dynamics, (22d) and (22e) are the trajectory and budget initial conditions based on the candidate trajectory and corresponding propagated budget state $b_{k}^{\textup{can},T_{S}}$, (22f) is the safety constraint, (22g) is the input constraint, and (22h) is the terminal constraint on the trajectory $p$.
We will solve (22) for various initial conditions with a sampling-based method detailed below.
@@ III-B2 ReRoot (Reverse Rooted Forest)
We build a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with a set of nodes $\mathcal{V}\subset\mathcal{X}$ and a set of edges $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$, in a similar manner to the standard RRT* algorithm [17]. The key differences between ReRoot and RRT* are as follows:
  • The graph $\mathcal{G}$ is a rooted forest, i.e., a union of disjoint rooted trees, where the roots are in the renewal set $\mathcal{R}_{k}$. Furthermore, in the rewiring step, it is possible for a node (and its children) to switch to a different tree with a different root.
  • We grow the forest in reverse such that the roots are backup “goal” states in $\mathcal{R}_{k}$ and the tree extends to cover $\mathcal{F}_{k}$, which results in optimal paths from all states in $\mathcal{F}_{k}$ that end in $\mathcal{R}_{k}$. In RRT* the initial state is the root and the tree is grown until it reaches the goal state.
  • Instead of minimizing path length, the paths minimize the budget state at the end of the backup trajectory $b^{\textup{bck}}$. Since the budget dynamics only depend on the state $x$ and control $u$, the cost of the path from a node to the root is the sum of the incremental costs along the path.
Despite these differences in the interpretation of the nodes and paths, the core RRT* algorithm is unchanged and provides asymptotic optimality of the cost of all paths as $|\mathcal{V}|\to\infty$.
Next, we describe the process of extracting a backup trajectory from the ReRoot graph.
@@ III-B3 Backup Trajectory Generation
Denote $d:\mathcal{V}\to\mathbb{R}_{\geq 0}$ as the depth of a node, which is the number of edges between $v$ and the root of its tree. A root node has depth $d=0$. For $v\in\mathcal{V}$, define
@@
as the set of waypoints from $v$ to the root of its tree, where $\mathcal{P}:\mathcal{V}\to\mathcal{V}$ is the parent node of $v$.
Then, for each $v\in\mathcal{V}$, let $p_{v}$ be the closed loop trajectory from tracking the waypoints $\mathcal{W}(v)$:
$\dot{p}_{v}=f(p_{v},\pi^{T}(\mathcal{W}(v)),~{}p_{v}(t_{kS})=p^{\textup{can},T_{S}}_{k}(t_{kS}).$ (24)
Additionally, let $b_{v}$ be the budget state expenditure over the trajectory $p_{v}$ by solving (3).
Finally, the trajectory with minimal $b_{v}$ that is safe for all time is chosen as the backup trajectory:
$p^{\textup{bck},T_{S}}_{k}=\operatorname*{\arg\!\min}_{p_{v}}\{b_{v}~{}|~{}p_{v}(t)\in\mathcal{F}_{k}~{}\forall t\geq t_{kS}\}.$ (25)
Note that by construction, $p^{\textup{bck},T_{S}}_{k}$ is dynamically feasible and guaranteed to remain in the known safe set $\mathcal{F}_{k}$. The resulting candidate trajectory $p^{\textup{can},T_{S}}_{k}$ is the solution to (11). The validity of the candidate trajectory is given from Definition 6, i.e., checking that $b(t)\leq B$ by forward propagating (3).
III-C Summary
In summary, we have provided a solution to 1 with a layered architecture built on the gatekeeper trajectory filter and the ReRoot backup policy. To guarantee that safety and budget constraints are satisfied for all time as the nominal trajectory is tracked, a backup trajectory is planned and stitched with the largest portion of the nominal trajectory such that the resulting committed trajectory satisfies all constraints and reaches a budget renewal set. We then provided a method for constructing backup trajectories called ReRoot that minimizes the expenditure of the budget state to allow the nominal trajectory to “consume” more of the budget. ReRoot is a reverse rooted forest in which branches represent paths of minimal budget cost that start from nodes in the known safe space and reach a budget renewal set.
@@ IV-A Simulation Setup
$\dot{r}_{N}$ $=V\cos\psi$ (26a)
$\dot{r}_{E}$ $=V\sin\psi$ (26b)
$\dot{\psi}$ $=u$ (26c)
where $[r_{N},r_{E}]^{\intercal}\in\mathbb{R}^{2}$ is the horizontal position in a North-East-Down (NED) coordinate frame, $\psi\in[0,2\pi)$ is the heading, $V>0$ is a fixed velocity and $u$ is the input turn rate. $u$ is constrained by the minimum turn radius of the UAV, $\rho>0$, such that $u\in\mathcal{U}=[-\frac{V}{\rho},\frac{V}{\rho}]$. For simplicity and visualization purposes, we constrain the UAV to fly at a fixed altitude, but our method extends trivially to 3D trajectories, or for more complicated dynamics.
The UAV’s velocity is $V=10$ m/s and minimum turn radius is $\rho=10$ m, which corresponds to a $\sim$45° bank angle. The sensor field of view (FoV) is defined by a radius of 60 m and an angle of 90°. The perception and mapping module runs at 5 Hz.
The objective to reach the randomly chosen goal location within the 250 m $\times$ 200 m mission domain. The mission requirements are as follows:
@@
Translating these requirements b into the format needed for our method, we first write the safe set as
@@
$\begin{cases}\dot{b}=0.03V,\quad&x\notin\mathcal{R}\\ b^{+}=b_{\textup{reset}}=0,\quad&x\in\mathcal{R}\end{cases}$ (28)
with initial condition
$b(t_{0})=\begin{cases}b_{0},\quad&x\notin\mathcal{R}\\ 0,\quad&x\in\mathcal{R}.\end{cases}$ (29)
@@
We initialize ReRoot around the starting location to form a tree of safe backups. Then, at 1 Hz, nodes are added based on the newly mapped features to extend the tree into the known safe set. As landmarks are discovered, a root node is placed and nearby nodes automatically connect to the new tree.
IV-B Results and Discussion
The Dubins dynamics in (26) allows fast backup trajectory generation because the edges of ReRoot represent dynamically feasible subpaths. Given two nodes $v_{1},v_{2}\in\mathcal{V}$, the Dubins path of minimum distance is unique and is found algebraically [22]. Hence any list of waypoints $\mathcal{W}(v)$ can be tracked perfectly and the budget state $b(t)$ along the backup trajectory is exactly the node cost of $v$, which is used when building ReRoot to determine the optimal parent of a new node. In the more general case, the trajectories can be computed numerically by solving the differential equations, as in [14].
@@ IV-B Results and Discussion
The code and animations of the simulation are available here [^1] https://github.com/dcherenson/gatekeeper_budget_constraints. In the simulation, we observe that the UAV reaches the goal while satisfying all constraints in Figure 4. Figure 4a shows that the nominal trajectory across the sparse field is found to be unsafe, so a short backup trajectory is planned to return to its orbit. In Figure 4b, the UAV has not mapped enough of the area around the mid-point landmark to reach it safely and must return to the starting location to reset the localization error. Then, in Figure 4c, the UAV reaches the landmark using the ReRoot tree rooted at the landmark. Finally, in Figure 4d, the UAV is able to reach the goal location with a committed trajectory back to the landmark. Figure 4e shows the position error over time which never exceeds the budget constraint. Figure 4f shows the number of visible features over time, showing that the safety constraint is always satisfied.
@@
We also simulate the mission with a lower, more challenging budget of $B=5$ m of accumulated localization error, which makes the task of reaching the goal while satisfying all constraints infeasible. In Figure 5a, the UAV is unable to reach the landmark or the goal without exceeding the limit. The resulting pattern is caused by the UAV following the nominal trajectory until gatekeeper prevents the UAV from continuing to avoid exceeding the budget, then returning to the home landmark, and the cycle continues. Figure 5b and Figure 5c show that the safety and budget constraints are satisfied, despite this infinite looping behavior. This demonstrates gatekeeper’s ability to ensure constraint satisfaction even when the mission objectives cannot be satisfied safely.
@@ V Conclusion
In this paper, we proposed an architecture centered on gatekeeper, an online safety verification algorithm, to guarantee that safety and budget constraints are enforced throughout a mission. We proved that the method is recursively feasible and guarantees safety for all time. We then introduced ReRoot, a method to construct a backup trajectory that allows for efficient online trajectory generation to minimize the expenditure of the budget state. The algorithm grows multiple rooted trees in reverse starting from budget renewal sets and extending into the known safe set.
We demonstrated our algorithm in simulation with a case study of a fixed-wing UAV in a GNSS-denied environment navigating to a goal location. The safety constraint was to maintain a minimum of 8 visual features in the FoV and the budget constraint was to limit the localization error to below 9 m. The combination of gatekeeper with ReRoot showed that multiple backup sets were used to allow the UAV to explore and map features while always staying in range of a renewal set to satisfy the budget constraint.

### v2

@@ 14 $T_{S}\leftarrow T_{S}-\Delta T$
Therefore, $p_{k}^{\textup{com}}(t)\in\mathcal{S}~\forall t\in[t_{k},t_{k+1}),~\forall k\in\mathbb{N}\implies p(t)\in\mathcal{S}~\forall t\geq t_{0}$. Moreover, the solution to (3) with the corresponding $p(t)$ and $u(t)$ satisfies $b(t)\leq B~\forall t\geq t_{0}$. ∎
We have shown the gatekeeper framework guarantees constraint satisfaction when paired with a suitable backup policy that ensures all candidate trajectories reach $\mathcal{R}_{k}$. An open challenge is how to construct a general backup policy that applies to a wide range of nonlinear system dynamics and non-convex safety constraints. To address this challenge, we introduce an algorithm to efficiently construct backup trajectories as the robot gathers information in the environment.
III-B ReRoot for Backup Trajectory Construction
We leverage the well-known RRT* path planner in a novel manner to propose a sampling-based algorithm called ReRoot that efficiently generates a dynamically feasible trajectory from any location to a backup set in which the budgeted resource is renewed.
III-B1 Backup Trajectory Generation Problem
@@
where $t_{kB}=t_{k}+T_{S}+T_{B}$ is the finite time when $p$ reaches $\mathcal{R}_{k}$. For any switch time $T_{S}$, the ideal backup trajectory is the one that minimizes the budget expenditure. The goal is to find such a backup trajectory to form the gatekeeper candidate trajectory.
Problem 2.
At the $k$-th iteration and for a given switch time $T_{S}$, given the known budget renewal set $\mathcal{R}_{k}$ at time $t_{k}$, first return a set of dynamically feasible trajectories starting at $p_{k}^{\textup{can},T_{S}}(t_{kS})$ and reaching $\mathcal{R}_{k}$ while remaining in the known safe set $\mathcal{F}_{k}$. Select among them the trajectory $p^{\textup{bck},T_{S}}_{k}$ with minimal budget expenditure $b^{-}$.
@@
2 for $k\in[1,\dotsc,K]$ do
3  $\mathcal{F}_{k},\mathcal{R}_{k}\leftarrow\textnormal{{UpdateFreeSpace}}()$
4  if $\mathcal{R}_{k}\cap\mathcal{R}_{k-1}\neq\emptyset$ then
5     $\textnormal{{AddRootNodes}}(\mathcal{G}_{k},\mathcal{R}_{k}\cap\mathcal{R}_{k-1})$
6
7 $\textnormal{{GrowReRoot}}(\mathcal{G}_{k},\mathcal{R}_{k},\mathcal{F}_{k},n_{\textup{update}})$
8  $p_{k}^{\textup{com}}=\textnormal{{gatekeeper}}(p_{k}^{\textup{nom}},\mathcal{G}_{k},x(t_{k}),b(t_{k}))$
9  Track $p_{k}^{\textup{com}}$
10  Update $b(t)$ via (3)
11
Algorithm 2 Main Autonomy Loop
@@ III-B2 ReRoot (Reverse Rooted Forest)
We build a graph $\mathcal{G}_{k}=(\mathcal{V}_{k},\mathcal{E}_{k})$ with a set of nodes $\mathcal{V}_{k}\subset\mathcal{X}$ and a set of edges $\mathcal{E}_{k}\subset\mathcal{V}_{k}\times\mathcal{V}_{k}$. The underlying algorithm for building $\mathcal{G}_{k}$ is either the standard RRT* algorithm [18], or extensions of RRT* that attempt to generate kinodynamically feasible paths [19, 20]. The novelty in our approach is the manner of initialization and interpretation of the generated graph $\mathcal{G}_{k}$. Algorithm 2 details the overall autonomy algorithm that runs at each iteration $k$, including the novel ReRoot initialization and growth steps. We summarize the key steps of ReRoot:
  • Root Nodes: The graph $\mathcal{G}_{k}$ is a rooted forest, i.e., a union of disjoint rooted trees where all root nodes are in the renewal set $\mathcal{R}_{k}$. The initialization step, Algorithm 2, creates the set of root nodes $\mathcal{V}_{0}$ in $\mathcal{R}_{0}$. During the mission, further updates via Algorithm 2 add more root nodes to $\mathcal{V}_{k}$ in the expanded renewal set $\mathcal{R}_{k}$. The root nodes serve as the endpoints of the backup trajectories.
  • Online Growth: In Algorithm 2 at each iteration $k$, the RRT* algorithm is used to add $n_{\rm update}$ nodes to $\mathcal{G}_{k}$, including the standard rewiring step. The path cost function used in RRT* is the budget dynamics (3) evaluated along the path. During the rewiring step, a node can switch trees, meaning that the path to the root of the new tree has a lower cost than the node’s path to its previous root node.
  • Path of Waypoints: From any node $v\in\mathcal{V}_{k}$ in the forest, there is a single path of waypoints $\mathcal{W}(v)$ to a root node in $\mathcal{R}_{k}$, which is not necessarily dynamically feasible. This path is used to generate a dynamically feasible trajectory by forward propagation of the closed-loop dynamics, which we describe below. This step occurs in Algorithm 1 of Alg. 1, called at each switching time $T_{S}$.
Remark 1.
Some popular variants of RRT/RRT* involve growing two trees with one root at the initial location and one at the goal location [21, 22, 23]. By contrast, ReRoot grows multiple trees by placing new roots as more components of the renewal set $\mathcal{R}_{k}$ are discovered online.
Next, we describe the in further detail the process of extracting a backup trajectory from the ReRoot graph during the gatekeeper candidate trajectory propagation step (Algorithm 1 of Algorithm 1).
@@ III-B3 Backup Trajectory Generation
Denote $d:\mathcal{V}_{k}\to\mathbb{R}_{\geq 0}$ as the depth of a node, which is the number of edges between $v$ and the root of its tree. A root node has depth $d=0$. For $v\in\mathcal{V}_{k}$, define
@@
as the set of waypoints from $v$ to the root of its tree, where $\mathcal{P}:\mathcal{V}_{k}\to\mathcal{V}_{k}$ is the parent node of $v$. Let $\mathcal{V}_{\rm near}$ be the set of all nodes within a ball of radius $R>0$ of the desired starting location of the backup trajectory, that is,
$\mathcal{V}_{\rm near}=\{v\in\mathcal{V}_{k}~|~\left\|v-p^{\textup{nom}}_{k}(t_{kS})\right\|\leq R\}.$ (20)
Then, for each $v\in\mathcal{V}_{\rm near}$, we generate the corresponding trajectory $(p_{v},u_{v})$ by propagating the closed loop trajectory of tracking the waypoints $\mathcal{W}(v)$ over the horizon $t_{kS},t_{kB}]$:
$\dot{p}_{v}(t)$ $=f(p_{v}(t),u_{v}(t)),~p_{v}(t_{kS})=p^{\textup{nom}}_{k}(t_{kS}),$ (21)
$u_{v}(t)$ $=\pi^{T}(p_{v}(t),\mathcal{W}(v))$ (22)
where $\pi^{T}:\mathcal{X}\times 2^{\mathcal{V}}_{k}\to\mathcal{U}$ is a waypoint tracking controller [^2] The path is not required to be tracked perfectly. Safety violations of the propagated trajectory is checked in the validation step., e.g. see [24, 25, 26]. $2^{\mathcal{V}}_{k}$ is the power set of $\mathcal{V}_{k}$, i.e., the set of all sets of $\mathcal{V}_{k}$. This forms a set $\mathcal{N}_{k}^{T_{S}}\subset\Phi(t_{kS},p_{k}^{\rm nom}(t_{kS}))$ of $|\mathcal{V}_{\rm near}|$ trajectories. Additionally, let $b_{v}^{-}$ be the budget state expenditure over the trajectory $p_{v}$ by solving (3). Finally, the trajectory with minimal $b_{v}^{-}$ that is safe for all time is selected as the backup trajectory:
$p^{\textup{bck},T_{S}}_{k}=\operatorname*{\arg\!\min}_{(p_{v},u_{v})\in\mathcal{N}_{k}^{T_{S}}}\{b_{v}^{-}~|~p_{v}(t)\in\mathcal{F}_{k},~\forall t\geq t_{kS}\}.$ (23)
The resulting candidate trajectory $p^{\textup{can},T_{S}}_{k}$ is validated by Definition 7, i.e., checking that $b^{-}\leq B$ by forward propagating (3) along $p^{\textup{can},T_{S}}_{k}$. If there does not exist a feasible backup trajectory for a given switching time $T_{S}$, then the gatekeeper algorithm decrements $T_{S}$ until either a feasible backup trajectory is found or $T_{S}=0$ and the previous committed trajectory $p_{k-1}^{\rm com}$ is selected as the next committed trajectory.
In summary, we solve 2 by selecting the backup trajectory of minimal budget expenditure among a set of potential backup paths in the ReRoot forest. Through the forward propagation in (21) and the validation step in (23), we show that the backup trajectory is dynamically feasible and safe by construction. Finally, the gatekeeper committed validation step checks budget constraint satisfaction of the candidate trajectory.
@@ IV Simulation Case Study
$\dot{r}_{N}=V\cos\psi,~\dot{r}_{E}=V\sin\psi,~\dot{\psi}=u$ (24)
where $r_{N},r_{E}]^{\top}\in\mathbb{R}^{2}$ is the horizontal position in a North-East-Down (NED) coordinate frame, $\psi\in[0,2\pi)$ is the heading, $V>0$ is a fixed velocity and $u$ is the input turn rate. $u$ is constrained by the minimum turn radius of the UAV, $\rho>0$, such that $u\in\mathcal{U}=[-\frac{V}{\rho},\frac{V}{\rho}]$. For simplicity and visualization purposes, we constrain the UAV to fly at a fixed altitude, but our method extends trivially to 3D trajectories, or for more complicated dynamics.
The UAV’s velocity is $V=10$ m/s and minimum turn radius is $\rho=10$ m, which corresponds to a $\sim$45° bank angle. The sensor field of view (FoV) is defined by a radius of 60 m and an angle of 90°. The perception and mapping module runs at 5 Hz. The objective is to reach the randomly chosen goal location within the 250 m $\times$ 200 m mission domain. The mission requirements are as follows:
@@
We write the safe set as
@@
$\begin{cases}\dot{b}=0.03V,\quad&x\notin\mathcal{R}\\ b^{+}=b_{\textup{reset}}=0,\quad&x\in\mathcal{R}.\end{cases}$ (26)
@@
We initialize ReRoot around the starting location to form a small tree of safe backups that return to the initial orbit. Then, at 1 Hz, nodes are added based on the newly mapped features to extend the tree into the known safe set. As landmarks are discovered, a root node is placed and nearby nodes automatically connect to the new tree.
The Dubins dynamics in (24) allows fast backup trajectory generation because the edges of ReRoot represent dynamically feasible subpaths. Given two nodes $v_{1},v_{2}\in\mathcal{V}_{k}$, the Dubins path of minimum distance is unique and is found algebraically [31]. The path of waypoints $\mathcal{W}(v)$ is dynamically feasible and the budget state $b(t)$ along the backup trajectory is the node cost of $v$, which is used when building ReRoot to determine the parent of a new node.
@@
The code and animations of the simulation are available here [^3] https://github.com/dcherenson/budget-constrained-planning. In the simulation, we observe that the UAV reaches the goal while satisfying all constraints in Figure 4. Figure 4a shows that the nominal trajectory across the sparse field is found to be unsafe, so a short backup trajectory is planned to return to its orbit. In Figure 4b, the UAV has not mapped enough of the area around the mid-point landmark to reach it safely and must return to the starting location to reset the localization error. Then, in Figure 4c, the UAV reaches the landmark using the ReRoot tree rooted at the landmark. Finally, in Figure 4d, the UAV is able to reach the goal location with a committed trajectory back to the landmark. Figure 4e shows the position error over time which never exceeds the budget constraint. Figure 4f shows the number of visible features over time, showing that the safety constraint is always satisfied. Table I shows the computation times of one iteration of Algorithms 2 and 2 in Algorithm 2 for ReRoot and one iteration of Algorithm 2 for gatekeeper.
TABLE I: Computation Times per Iteration
Component Mean [ms] Std. Dev. [ms]
ReRoot 2.90 7.17
gatekeeper 2.52 4.48
@@
We also simulate the mission with a lower, more challenging budget of $B=5$ m of accumulated localization error, which makes the task of reaching the goal while satisfying all constraints infeasible. In Figure 5a, the UAV is unable to reach the landmark or the goal without exceeding the limit. The resulting pattern is caused by the UAV following the nominal trajectory until gatekeeper prevents the UAV from continuing to avoid exceeding the budget, then returning to the home landmark, and the cycle continues. Figure 5b and Figure 5c show that the safety and budget constraints are satisfied, despite this infinite looping behavior, which demonstrates the guaranteed constraint satisfaction.
@@ V Conclusion
In this paper, we proposed an architecture to guarantee safety and budget constraints throughout a mission in an environment where the safe set is built on-the-fly. The key contribution is ReRoot, a sampling-based backup planning framework that augments the gatekeeper architecture to construct backup policies, which are needed for guaranteeing constraint satisfaction. By growing multiple reverse RRT* trees rooted in renewal sets, ReRoot efficiently generates trajectories that satisfy constraints while minimizing resource expenditure. The efficacy of our approach was demonstrated in simulation with a case study of a fixed-wing UAV in a GNSS-denied environment navigating to a goal location. The safety constraint was to maintain a minimum of 8 visual features in the FoV and the budget constraint was to limit the localization error to below 9 m. gatekeeper with ReRoot allowed the UAV to explore and map features while always satisfying the budget and safety constraints.